OpenStudy (anonymous):

what is 1/8 + 3/4 = 1/8 + ?/8 = what?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

1/8+3/4=1/8+6/8=7/8

OpenStudy (anonymous):

6 and 7/8

OpenStudy (anonymous):

thanks